Circular No. 8453

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004gs IN MCG +03-22-20 Further to IAUC 8452, S. Park and W. Li report the LOSS discovery

of an apparent supernova (mag 17.1) on an unfiltered KAIT image taken on Dec. 12.43 UT. SN 2004gs is located at α = 8^{h}38^{m}23^{s}.18, δ = +17^{o}37′39′′.8 (equinox 2000.0), which is 9′′.8 west and 12′′.7 south of the nucleus of MCG +03-22-20. A KAIT image taken on Dec. 4.42 already showed a hint of the new object (limiting mag ∼ 19.5), while an image taken on Nov. 24.45 showed nothing at this position (limiting mag ∼ 19.0).

N. Morrell, G. Folatelli, and M. Hamuy, Carnegie Supernova Project, report that a CCD spectrogram (380–920 nm) of SN 2004gs, obtained on Dec. 13.25 UT with the Las Campanas 2.5-m du Pont telescope (+ WFCCD Spectrograph), shows it to be a type-Ia supernova near maximum light. The expansion velocity, derived from the blueshift of the minimum of Si II 635.5-nm, is 11600 km/s (assuming a recession velocity of 7988 km/s for the host galaxy as given in the NED database). SN 2004gs shows an enhanced Si II 597.2-nm absorption line. No evidence for Na I D 589.3-nm can be seen (with an upper limit of 0.05 nm), which suggests little or no dust absorption in the host galaxy.

SUPERNOVAE 2004gd, 2004ge, 2004gf, AND 2004gl Further to IAUC 8452, A. V. Filippenko and R. J. Foley write that

inspection of CCD Keck I spectra, also obtained on Dec. 12 UT, shows that SN 2004gd (IAUC 8443) is of type IIn, strongly resembling SN 1985G (Pastorello et al. 2002, MNRAS 333, 27). Prominent, relatively narrow (FWHM 1000 km/s) hydrogen-Balmer emission lines are superposed on broader bases and exhibit narrow P-Cyg absorption components with minima displaced by ∼ 700 km/s from the emission-line cores. Similar absorption features are seen in the Fe II multiplets and other lines; they are probably produced by a dense circumstellar shell. SN 2004ge (IAUC 8443) is of type Ic, shortly after maximum brightness; the shape of its continuum argues that it is heavily reddened, and the equivalent width of the narrow interstellar Na I D absorption line at 589 nm is ∼ 0.2 nm. SN 2004gf (IAUC 8444) is of type II, probably within a few weeks past explosion, based on the blue continuum; the Hα line is vastly dominated by the emission component (FWHM = 2200 km/s), although all other lines (Balmer, Fe II, O I, Ca II) show more typical P-Cyg profiles. SN 2004gl (IAUC 8446) is of type Ia, three weeks past maximum brightness; the redshift of the host galaxy, measured from narrow emission lines in the nucleus, is 0.0393.

2004 December 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green